What are the steps to turn this equation 1.50x+1.25y=375 into an equation you can graph?
so far i have
1.50x+1.25y=375
-1.50x -1.50x
1.25y=375-1.50x
and now i dont know what else to do or if im even on the right track.

Answer by RAY100(1637)

Very good so far.
Now,, we make a T box,,and just plot some "arbitrary" points
.
x,,,,,,,,,,,,y
.
0,,,,,,,,,,,,300,,,,,,,,,,y= 375 / 1.25 = 300,,,assume x=0,,calc y
.
250,,,,,,,,,,,0,,,,,,,,,,,x=375/1.5=250,,,assume y=0,,,calc x
.
125,,,,,,,,,,150,,,,,,,,,,1.5(125) +1.25y = 375,,y=(375-187.5)/1.25 =150
.
just plot the coordinates on the x-y grid, and connect the dots.
.
have a good day